OpenStudy (anonymous):

Derivative of ytan(x+y)=4 ?

myininaya (myininaya):

derivative of 4 is 0 for the other side its a bit trickier you need to recall the product rule and also the chain rule \[[y(\tan(x+y)]'=y'\tan(x+y)+y[\tan(x+y)]'\] now need to find \[[\tan(x+y)]'\] which needs chain rule \[[\tan(x+y)]'=(x+y)'\sec^2(x+y)=(1+y')\sec^2(x+y)\]

OpenStudy (anonymous):

\[\frac{dy}{dx}\times \tan(x+y) +y \times \sec^2{(x+y)} \left( 1 + \frac{dy}{dx}\right)\]

OpenStudy (anonymous):

oh wow

myininaya (myininaya):

so you have \[y' \tan(x+y)+y (1+y')\sec^2(x+y)=0\]

OpenStudy (anonymous):

Thanks so much- any way to get dy/dx on one side by itself?

myininaya (myininaya):

yes thats what you do is solve for y'

myininaya (myininaya):

do some distribution put your terms that have y' together and everything without the factor y' on the other side then factor the y' out of the terms that did have y' in common and divide on both sides by whatever is being multiplied by y'

OpenStudy (anonymous):

So after distributing:\[dy/dx * \tan (x+y) + y(\sec ^{2}(x+y)+ dy/dx * \sec ^{2}(x+y)\] Correct?

OpenStudy (anonymous):

Pre- solving, post- distribution.

myininaya (myininaya):

ok so the close parenthesis is meant to be at the end if so then so far good also distribute y to both of those terms inside that parenthesis ( and i think you meant to close at the very end)

OpenStudy (anonymous):

\[dy/dx∗\tan(x+y)+(ysec2(x+y)+dy/dx∗ysec2(x+y))\] This? Then solve for dy/dx?

OpenStudy (anonymous):

Those are sec^2, not sec 2(x+y).

myininaya (myininaya):

ok good

myininaya (myininaya):

now put your terms that have y' together and put everything else on the other side of equation

OpenStudy (anonymous):

\[\left(\begin{matrix}dy \\ dx\end{matrix}\right) \tan (x+y) + \left(\begin{matrix}dy \\ dx\end{matrix}\right) y \sec ^{2}(x+y) = -y \sec ^{2}(x+y)\] Factor out dy/dx and divide by what is then (tan(x+y) + y sec^2(x+y)), leaving \[\left(\begin{matrix}dy \\ dx\end{matrix}\right) = \left(\begin{matrix}(-y \sec ^{2}(x+y)) \\ (\tan (x+y) + y \sec ^{2}(x+y))\end{matrix}\right)\] Is that the answer?

myininaya (myininaya):

looks awesome

OpenStudy (anonymous):

Woohoo! Thank you so much, I really appreciate it.